# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v2

@@ Appendix C Functional inequalities
[Che70] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^5] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for log-concave distributions due to [Mil09], while $C_{\mathsf{PI}}(\mu)\gtrsim C_{\mathsf{Ch}}^{-2}(\mu)/d$ for general distributions due to [Led94]., and then [KLS95] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\lVert X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\rVert^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}\Sigma_{\mu_{\mathsf{LC}}})^{1/2}}\,.$ (C.1)
@@
Kannan et al. proposed the KLS conjecture in the same paper, which says that for the spectral norm $\lVert\cdot\rVert_{2}$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{1}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,.$
While the original result in [KLS95] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic log-concave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case. Following a line of work [LV17, Che21, KL22, Kla23], the current bound is
$C_{\mathsf{Ch}}(\mu_{\mathsf{LS}})\gtrsim\frac{(\log d)^{-1/2}}{\lVert\Sigma_{\mu_{\mathsf{LS}}}\rVert_{2}^{1/2}}\,,$
@@
[Led94] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ and $C_{\mathsf{LSI}}(\mu_{\mathsf{LS}})\gtrsim C_{\mathsf{logCh}}^{-2}(\mu_{\mathsf{LS}})$. [KLM06] showed that any log-concave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1}$. Later in 2016, [LV17] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LS}})\gtrsim D^{-1/2}$ under isotropy. Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
Appendix D The Wasserstein geometry
We present additional technical background on the Wasserstein geometry and Markov semigroup theory. Interested readers can refer to [Vil09, AGS05, Che24] for standard references on Wasserstein spaces and applications to sampling.
Wasserstein gradient.
Let $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ be the space of probability measures admitting densities on $\mathbb{R}^{d}$ with finite second moment. Although there are many ways to metrize $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, the geometry induced by the Wasserstein-$2$ distance $\mathcal{W}_{2}$ is a particularly useful structure for analysis.
Under the $\mathcal{W}_{2}$-geometry, one can define a “gradient” of a functional defined over $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$. Specifically, for a functional $\mathcal{F}:\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})\to\mathbb{R}\cup\{\infty\}$, the Wasserstein gradient of $\mathcal{F}$ at $\mu\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ is defined as $\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu)=\nabla(\delta\mathcal{F})(\mu)\in L^{2}(\mu)$, where $\nabla$ is the standard gradient and $\delta\mathcal{F}$ is the first variation of $\mathcal{F}$ [^6] The first variation can be defined, for any measures $\nu_{0},\nu_{1}\in\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$, as $\lim_{t\downarrow 0}\frac{\mathcal{F}((1-t)\nu_{0}+t\nu_{1})-\mathcal{F}(\nu_{0})}{t}=\langle\delta\mathcal{F}(\nu_{0}),\nu_{1}-\nu_{0}\rangle$. This definition is unique up to an additive constant, which is irrelevant as we are only concerned with its gradient.. Equipped with this $\mathcal{W}_{2}$-gradient, one can define the Wasserstein gradient flow of $\mathcal{F}$ that describes the evolution of a measure $\{\mu_{t}\}_{t\geq 0}$, from some initial measure $\mu_{0}$, as follows:
$\partial_{t}\mu_{t}=\operatorname{div}\bigl{(}\mu_{t}\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu_{t})\bigr{)}\,.$
More generally, we can identify the Wasserstein “velocity” for some measure $\mu_{t}$ as $v_{t}$ if the time derivative of $\mu_{t}$ can be written in the form
$\partial_{t}\mu_{t}=-\operatorname{div}(\mu_{t}v_{t})\,.$
Under this identification, the time derivative of a functional $\mathcal{F}$ on $\mathcal{P}_{2,\rm ac}(\mathbb{R}^{d})$ with smooth Wasserstein gradient under these dynamics can be written as
$\partial_{t}\mathcal{F}(\mu_{t})=\mathbb{E}_{\mu_{t}}\langle\nabla_{\mathcal{W}_{2}}\mathcal{F}(\mu_{t}),v_{t}\rangle\,,$
when $v_{t}\in\overline{\{\nabla\psi:\psi\in C_{c}^{\infty}(\mathbb{R}^{d})\}}^{L^{2}(\mu_{t})}$, where {}¯$\overline{\{\cdot\}}^{L^{2}(\mu_{t})}$ denotes the closure of a set with respect to $L^{2}(\mu_{t})$. This is the appropriate notion of tangent space in this geometry.
For instance, when we take the functional to be the entropy of the measure, $\mathcal{H}(\mu)\coloneqq\frac{1}{2}\int\mu\log\mu$, one can verify $\nabla_{\mathcal{W}_{2}}\mathcal{H}(\mu)=\frac{1}{2}\nabla\log\mu$. The heat flow equation can be written as $\partial_{t}\mu_{t}=\frac{1}{2}\Delta\mu_{t}=\frac{1}{2}\operatorname{div}(\nabla\mu_{t})=\frac{1}{2}\operatorname{div}(\mu_{t}\nabla\log\mu_{t})$, which indicates that the velocity of measures $\mu_{t}$ under the heat flow is $v_{t}=-\frac{1}{2}\nabla\log\mu_{t}$. Hence, we can notice that $\nabla_{\mathcal{W}_{2}}\mathcal{H}(\mu_{t})=-v_{t}$, and thus recover the heat flow as the Wasserstein gradient flow of the entropy of the measure.
Fokker-Planck equation and time-reversal of SDE.
Consider a stochastic differential equation $(X_{t})$ given by
$\mathrm{d}X_{t}=-a_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}\,.$ (D.1)
@@ Appendix D The Wasserstein geometry
$\partial_{t}\mu_{t}=\mathsf{div}(\mu_{t}a_{t})+\frac{1}{2}\Delta\mu_{t}\,,$ (D.2)
correspond to $\mathsf{law}(X_{t})$. In this context, (D.2) is referred to as the Fokker-Planck equation corresponding to (D.1).
From this equation, one can deduce the Fokker-Planck equation of the time reversal $\mu_{t}^{\leftarrow}:=\mu_{T-t}$:
@@
$\mathrm{d}X_{t}=\bigl{(}a_{T-t}(X_{t})+\nabla\log\mu_{T-t}(X_{t})\bigr{)}\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}^{\leftarrow}=\mu_{T}\,.$ (D.3)
While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot|z)$ if one starts (D.3) at $X_{0}=z$. This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform. The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [KP21] as a reference to its application in this context.

### v3

@@ Appendix B Functional inequalities
[Che70] established $C_{\mathsf{PI}}(\mu)\lesssim C_{\mathsf{Ch}}^{-2}(\mu)$ [^4] The opposite direction $C_{\mathsf{PI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{Ch}}^{-2}(\mu_{\mathsf{LC}})$ also holds for logconcave distributions due to [Led94]., and then [KLS95] showed that for covariance matrix $\Sigma_{\mu}:=\mathbb{E}_{\mu}[(\cdot-\mathbb{E}_{\mu}X)(\cdot-\mathbb{E}_{\mu}X)^{\mathsf{T}}]$,
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{(\mathbb{E}_{\mu_{\mathsf{LC}}}[\lvert X-\mathbb{E}_{\mu_{\mathsf{LC}}}X\rvert^{2}])^{1/2}}=\frac{1}{(\operatorname{tr}\Sigma_{\mu_{\mathsf{LC}}})^{1/2}}\,.$ (B.1)
@@
Kannan et al. proposed the KLS conjecture in the same paper, which says that
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{1}{\lVert\Sigma_{\mu_{\mathsf{LC}}}\rVert_{\mathsf{op}}^{1/2}}\,.$
While the original result in [KLS95] ensures $C_{\mathsf{Ch}}\gtrsim d^{-1/2}$ for an isotropic logconcave distribution (due to $\Sigma=I_{d}$), this conjecture indeed claims $C_{\mathsf{Ch}}\gtrsim 1$ for such case. Following a line of work [LV24, Che21, KL22, Kla23], the current bound is
$C_{\mathsf{Ch}}(\mu_{\mathsf{LC}})\gtrsim\frac{(\log d)^{-1/2}}{\lVert\Sigma_{\mu_{\mathsf{LC}}}\rVert_{\mathsf{op}}^{1/2}}\,$,
@@
[Led94] established that $C_{\mathsf{LSI}}(\mu)\lesssim C_{\mathsf{logCh}}^{-2}(\mu)$ and $C_{\mathsf{LSI}}(\mu_{\mathsf{LC}})\gtrsim C_{\mathsf{logCh}}^{-2}(\mu_{\mathsf{LC}})$. [FK99] showed that any logconcave distributions with support of diameter $D>0$ satisfy $C_{\mathsf{logCh}}(\mu_{\mathsf{LC}})\gtrsim D^{-1}$. Later in 2016, [LV24] improved this to $C_{\mathsf{logCh}}(\mu_{\mathsf{LC}})\gtrsim D^{-1/2}$ under isotropy. Therefore, for convex $\mathcal{K}$, it follows that $C_{\mathsf{LSI}}(\pi)\lesssim D^{2}$ and that $C_{\mathsf{LSI}}(\pi)\lesssim D$ if $\pi$ is isotropic.
Appendix C Semigroup calculus for contraction under the heat flow
In this part, we first recall Markov semigroups, which are well-established mathematical tools that abstract an underlying Markov process. Interested readers can refer to [BGL14, Che25] for references. We then review the contraction results for heat flow and its time-reversal [CCSW22], which are intimately connected with our algorithm. We also provide key technical ingredients needed for its proof, such as the computations for measures evolving under simultaneous forward/backward heat flows.
Forward heat flow.
We begin by introducing the “heat flow” equation (or also known as the Fokker–Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
$\mathbb{E}_{\mu_{h}}[f(x)]=\mathbb{E}_{\mu_{0}}[P_{h}f(x)]\,$,
[…]
Backward heat flow.
Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP23], which is the most immediately useful.
Given some initial measure $\nu$ and some time horizon $h$, the adjoint corresponds to reversing ($\mathsf{FH}$) for times in $[0,h]$ when the initial distribution under consideration is $Z_{0}\sim\nu$. For other measures, it must be interpreted more carefully, and is given by the following partial differential equation starting from some measure $\mu_{0}^{\leftarrow}$ (see (C.1) and its derivation):
The ultimate purpose of this machinery is to affirm our earlier description of the Gibbs sampling procedure as alternating forward and backward heat flows. Indeed, notice that, if $\mu_{i}^{X}$ is the law of the iterate at some iteration $i$, then $\mu_{i}^{X}P_{h}$ is precisely $\mu_{i+1}^{Y}$ under our scheme, while $(\mu_{i}^{X}P_{h})Q_{h}^{\pi^{X},h}$ is $\mu_{i+1}^{X}$, assuming $Q_{h}^{\pi^{X},h}$ is well defined for non-smooth measures $\pi^{X}$. Thus, while Algorithm 1 is implemented via discrete steps, it can be exactly analyzed through arguments in continuous time.
Fokker–Planck equation and time-reversal of SDE.  The description above can be further generalized further as follows. Consider a stochastic differential equation $(X_{t})$ given by
$\mathrm{d}X_{t}=-a_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}\,.$ (C.1)
@@ Appendix C Semigroup calculus for contraction under the heat flow
$\partial_{t}\mu_{t}=\mathsf{div}(\mu_{t}a_{t})+\frac{1}{2}\Delta\mu_{t}\,$, (C.2)
correspond to $\mathsf{law}(X_{t})$. In this context, (C.2) is referred to as the Fokker–Planck equation corresponding to (C.1).
From this equation, one can deduce the Fokker–Planck equation of the time reversal $\mu_{t}^{\leftarrow}:=\mu_{T-t}$:
@@
$\mathrm{d}X_{t}=\bigl(a_{T-t}(X_{t})+\nabla\log\mu_{T-t}(X_{t})\bigr)\,\mathrm{d}t+\,\mathrm{d}B_{t}\qquad\text{with }X_{0}\sim\mu_{0}^{\leftarrow}=\mu_{T}\,.$ (C.3)
While the law of this process will give $\mu_{T}^{\leftarrow}=\mu_{0}$ at time $T$, it is also true that it will give $\mu_{0|T}(\cdot\,|\,z)$ if one starts (C.3) at $X_{0}=z$. This is a subtle fact, whose justification requires the introduction of a tool called Doob’s $h$-transform. The presentation of this subject is beyond the scope of this paper, and we refer interested readers to [KP23] as a reference to its application in this context.
Contraction under simultaneous evolution.  Instead of considering the change in metrics along the evolution of $\mu P_{t}$ with respect to “fixed” $\nu$, it will be useful to consider the simultaneous evolution of $\mu P_{t},\nu P_{t}$ (and similarly $\mu Q_{t}^{\pi^{X},h},(\nu P_{h})Q_{t}^{\pi^{X},h}$). This type of computation was carried out for specific metrics in earlier work [VW23, CCSW22]. The following is a more generalized form of one appearing in [YFL+23, Lemma 2]. In the lemma below, we consider an arbitrary diffusion equation with corresponding Fokker–Planck equation:
$\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\,\mathrm{d}B_{t}\,\quad\text{and}\quad\partial_{t}\mu_{t}=-\nabla\cdot(b_{t}\mu_{t})+\frac{1}{2}\,\Delta\mu_{t}\,$ (C.4)
Below, we derive contractions of two measures along the same stochastic process, which proves Proposition 1 under suitable regularity assumptions.
Lemma 28 (Decay along forward/backward heat flows).
Let $(\mu_{t})_{t\geq 0},(\nu_{t})_{t\geq 0}$ denote the laws of the solutions to (C.4) starting at $\mu_{0},\nu_{0}$ respectively. Assume that $\mu_{t}$ and $\nu_{t}$ has sufficient regularity and fast tail decay so that we can discard boundary terms in integration by parts and differentiate under the integral sign. Then, for any differentiable function $g$,
Proof.
For brevity, we drop the variable $x$ from functions involved and proceed by differentiating under the integral sign and discarding boundary terms in integration by parts:
$\partial_{t}D_{f}(\mu_{t}\mathbin{\|}\nu_{t})$ $=\int\Bigl\{\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\partial_{t}\nu_{t}+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\bigl(\frac{\mu_{t}}{\nu_{t}}\bigr)^{\prime}\nu_{t}\Bigr\}\,\mathrm{d}x$
$=\int\Bigl\{\partial_{t}\nu_{t}\Bigl(\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)-\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\Bigr)+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\partial_{t}\mu_{t}\Bigr\}\,\mathrm{d}x$
$\underset{(i)}{=}\int\bigl[-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\,\Delta\nu_{t}\bigr]\Bigl(\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)-\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\Bigr)\,\mathrm{d}x$
where in $(i)$ we substitute the F-P equation from (C.4). Integrating by parts (i.e., $\int f\operatorname{div}(\mathbf{G})=-\int\langle\nabla f,\mathbf{G}\rangle$ for a real-valued function $f$ and vector-valued function $\mathbf{G}$), we have that
$\int\bigl[-\nabla\cdot(b_{t}\nu_{t})\bigr]\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\,\mathrm{d}x=\int\bigl\langle b_{t}\nu_{t},\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,.$ (C.5)
On the other hand, we have that
$\int\bigl[-\nabla\cdot(b_{t}\nu_{t})\bigr]\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=-\int\bigl\langle b_{t}\nu_{t},\frac{\mu_{t}}{\nu_{t}}\,\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,.$
The second term cancels with the RHS of (C.5). We have a similar cancellation for the $\frac{1}{2}\,\Delta\nu_{t}$ term:
and
$\int\frac{1}{2}\,\Delta\nu_{t}\,\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\,\mathrm{d}x=\int\frac{1}{2}\,\bigl\langle\nabla\nu_{t},\frac{\mu_{t}}{\nu_{t}}\,\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)+\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\nabla\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x\,.$
Combining these, we are left with
$\int\bigl[-\nabla\cdot(b_{t}\nu_{t})+\frac{1}{2}\,\Delta\nu_{t}\bigr]\Bigl(\bigl(f\circ\frac{\mu_{t}}{\nu_{t}}\bigr)-\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\Bigr)\,\mathrm{d}x$ $=-\int\bigl\langle b_{t}\nu_{t}-\frac{1}{2}\,\nabla\nu_{t},\nabla\bigl(f^{\prime}\circ\frac{\mu_{t}}{\nu_{t}}\bigr)\frac{\mu_{t}}{\nu_{t}}\bigr\rangle\,\mathrm{d}x$
Finally, we note that
which completes the proof. ∎